OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (texaschic101):

distribute through the parenthesis and what do you get ?

OpenStudy (anonymous):

-2x + -6 = (-2x + -2) -4

OpenStudy (texaschic101):

-2(x + 3) = -2(x + 1) + 4 -2x - 6 = -2x - 2 + 4 combine like terms -2x - 6 = -2x + 2 still with me ?

OpenStudy (anonymous):

Yes

OpenStudy (texaschic101):

-2x - 6 = -2x + 2 (add 2x to both sides) -2x + 2x - 6 = 2 (now add 6 to both sides) -2x + 2x = 2 + 6 0 = 8 understand so far ?

OpenStudy (anonymous):

Yes I do.

OpenStudy (texaschic101):

and since 0 does not equal 8, there is NO SOLUTION to this problem. IF it would have ended up like : 0 = 0 or 4 = 4 or something that is equal, then there would have been infinite solutions

OpenStudy (anonymous):

Okay, thank you TexasChic101

OpenStudy (texaschic101):

anytime :)